Two skaters, a man and a woman, are standing on ice. Neglect any friction between the skate blades and the ice. The mass of the man is 87 kg, and the mass of the woman is 49 kg. The woman pushes on the man with a force of 56 N due east. Determine the acceleration (magnitude and direction) of (a) the man and (b) the woman.
1
Expert's answer
2013-10-01T10:51:27-0400
1. Acceleration of the man is
a=F/m_m =56 N / (87kg) = 0.64 m/s^2
direction - due east

2. Acceleration of the woman is
a=F/m_w =56 N / (49kg) = 1.14 m/s^2
direction - due west